You are conducting a significance test of H0: μ = 5 against Ha: μ > 5. After checking the conditions are met from a simple random sample of 30 observations, you obtain t = 2.35. Based on this result, describe the p-value.

The p-value falls between 0.15 and 0.2.

The p-value falls between 0.025 and 0.05.

The p-value falls between 0.01 and 0.02.

The p-value falls between 0.005 and 0.01.

The p-value is less than 0.005.
Answer #1

Solution :

Given that,

This is the right tailed test ,

t = 2.35

n = 30

degrees of freedom = n - 1 = 30 - 1 = 29

Using t table,

0.01 < P-value < 0.02

The p-value falls between 0.01 and 0.02.